# Measuring AI Ability to Complete Long Tasks

### v2

@@ 4.1 From raw data to time horizon
The task-specific threshold is chosen to represent human performance. For HCAST, the task-specific threshold is the same “target score” the human baseliner tries to achieve, which we also use to filter for successful runs. RE-Bench tasks have a fixed time rating of 8 hours, so the task-specific threshold is the average score of 7-9–hour human runs.
@@
Once we have agent success rates and time ratings for each task, our approach is inspired by item response theory (IRT) [32]. Like in IRT, we use logistic regression to find the task difficulty at which the agent has a 50% chance of success, but unlike IRT, we exploit the fact that we have human baselines to use difficulty ratings based on human time rather than ratings learned from agent performance. Specifically, we perform logistic regression using:
$p_{\mathrm{success}}(\mathrm{model},\mathrm{task})=\sigma((\log h_{\mathrm{model}}-\log t_{\mathrm{task}})\cdot\beta_{\mathrm{model}})$
@@
where $t_{\mathrm{task}}$ is the geometric mean time of successful human baselines, and $h_{\mathrm{model}}$ and $\beta_{\mathrm{model}}$ are learned parameters, with $h_{\mathrm{model}}$ representing the 50% horizon time. Further details and comparison to standard IRT methods are provided in Appendix B.4.
Excess success rates
Excess success rates ($\frac{S_{observed}-S_{predicted}}{S_{predicted}}$) are a metric for how much better (or worse) an AI agent performed compared to what we would expect, given a task’s length and that model’s ability (expected success can be seen in Figure 5). The average correlation between model excess success rates across all tasks is 0.40, indicating that AI agents still have moderately correlated performance when controlling for task length (See Figure 23 for the full correlation matrix).
4.2 Model horizon length vs. release date
[…]
@@
Horizon length on our tasks increases substantially over the entire time period from 2019 to early 2025. Base models like GPT-3 can complete some tasks requiring only writing a few words, but fail all tasks above 1 minute. Chat models like GPT-4 and Claude 3 Opus are able to complete the easier HCAST tasks with some frequency, leading to time horizons in the 5–30 minute range. However, the trend in 2024 and early 2025 may be faster, with o1 and Claude 3.7 Sonnet lying above the long-run trend. Though the gap is difficult to distinguish from noise, it is robust to methodological ablations like using continuous scoring (Appendix D).
@@ 4.2.1 Time horizons at 50% success rate vs 80% success rate
To check whether our choice of 50% success rate affects the long-run trend, we also compute the time horizon at which AI agents succeed at tasks with 80% success rate, shown in Figure 6. The doubling time in 80% time horizon (213 days) is similar to the doubling time of 50% time horizon (212 days), within margin of error. However, there is a large gap between models’ 50% time horizon and 80% time horizon—Claude 3.7 Sonnet has the longest 80%-horizon among models we examined at around 15 minutes, in contrast to its 50%-horizon of 59 minutes. This gap suggests that even models that sometimes succeed on difficult and diverse tasks cannot reliably perform tasks of moderate length. Due to our limited dataset, we cannot confidently measure time horizons at very high success rates (e.g. 95%)– see Section 8.1.
5 Qualitative analysis
To better contextualize the observed trend of improved model performance, we examined the transcripts for the tasks where earlier models (e.g. GPT-4) do substantially worse than current models. Specifically, we categorized our task families based on the type of expertise required, and noticed that current models tend to outperform earlier models for tasks that involved ML training, reverse engineering compiled software binaries, and cybersecurity CTFs. In addition, we noticed that agents improved greatly on tasks requiring situational awareness of the AI agent’s limitations or defeating an opposing strategy. This gave us five sets of task families to examine.
For each of these task family sets, we then used contractors to manually read through all runs from all models for all tasks in each of these five task families, and identified possible explanations for the improvement in AI agent performance on those tasks, as well as potential limitations. We find that models seem to have improved greatly in terms of tool use capabilities, demonstrate a markedly greater ability to adapt to mistakes (as opposed to repeating unsuccessful actions), and perform much better at parts of tasks requiring logical reasoning or code generation. However, we noticed that AI agents still seem to struggle in intuitively “messier” environments—specifically, environments without clear feedback loops, or where the agent needs to proactively seek out relevant information. We provide examples of both the improvements and these limitations in Appendix C.
We report the results in Table 3. We find that over a third of the GPT-4 failures resulted from repeating failed actions, compared to 2 out of 32 for o1, which we see as quantitative evidence for our claim that models seem to have improved in their ability to adapt to mistakes. Interestingly, half of the o1 failures resulted from abandoning the task prematurely, while only a quarter of the GPT-4 failures resulted from the same—this may result from o1 failures occurring on qualitatively more difficult tasks, or may reflect idiosyncrasies of o1.
Failure type GPT-4 1106 o1
@@ 5 Qualitative analysis
Table 3: Number of different categories of failures for 31 failed runs by GPT-4 1106 and 32 failed runs by o1 (Section 5). Note that as o1 succeeds at more tasks, its failures correspond to more challenging tasks compared to GPT-4’s failures.
@@ 6 External validity and robustness
To investigate the applicability of our results to other benchmarks, and to real task distributions, we performed four supplementary experiments. First, we check whether the 2023–2025 trend without the SWAA dataset retrodicts the trend since 2019, and find that the trends agree. Second, we label each of our tasks on 16 “messiness” factors—factors that we expect to (1) be representative of how real-world tasks may systematically differ from our tasks and (2) be relevant to AI agent performance. Third, we calculate AI agent horizon lengths from SWE-bench Verified tasks. We find a similar exponential trend, although with a shorter doubling time. However, we believe this shorter doubling time to be a result of SWE-bench Verified time annotations differentially underestimating the difficulty easier SWE-bench tasks. Finally, we collect and baseline a small set of uncontaminated issues from internal METR repositories. We find that our contracted human baseliners take much longer to complete these tasks than repository maintainers. We also find that AI agent performance is worse than would be predicted by maintainer time-to-complete but is consistent with contractor time-to-complete, given the AI agent success curves from HCAST + SWAA + RE-Bench tasks shown in Figure 5.
6.1 Retrodiction from 2023–2025 data
As part of exploratory work for this paper, we measured the time horizon of 9 frontier and near-frontier models released in 2023 and 2024, using only our HCAST and RE-Bench suites. This trend (with Claude 3.7 Sonnet added) is shown in Figure 7: time horizon doubles about every six months. Since we only had two 2023 models (GPT-4 0314 and GPT-4 1106) and a small data range (release date spanning 2 years and time horizon spanning 5 doublings), error bars were very wide. In addition, restricting our data further to 2024-only models produced a different trend with time horizon doubling about every three months, so any extrapolation into the future would not be robust.
To address these issues, we collected more data to extend the trendline into the past, developing the Software Atomic Actions (SWAA) suite to decrease the minimum human time of our task suite from 1 minute to under 2 seconds, and enabling us to measure GPT-2, davinci-002 (GPT-3) and GPT-3.5-turbo-instruct on the combined suite.
The 2023–2025 trend retrodicts the longer-term trend well (Figure 8). The measured doubling time over the whole 6 year period 2019–2025 inclusive was 212 days, which closely matches the 191-day trend based on data from non-SWAA tasks and 2023–2025 models.
6.2 Messiness factors
Real-world intellectual labor often involves messy details that benchmarks usually don’t include, such as being under-specified or poorly scoped, having unclear feedback loops or success criteria, or requiring coordination between multiple streams of work in real-time. We generally observed that agents struggle more on tasks that have these “messy” details (Section 5). A natural question is therefore whether agents showed similar rates of improvement on “less messy” and “more messy” tasks.
Figure 9: Performance trends over time for HCAST and RE-Bench tasks by length and messiness (Section 6.2). The data spans only 2023–2024 as pre-2023 models score 0 on non-SWAA tasks. Whilst our messier tasks have lower average success rates, trends in model performance improvements are not obviously slower on the high messiness split.
We rated HCAST and RE-Bench tasks on 16 properties that we expected to be 1) representative of how real world tasks might be systematically harder than our tasks and 2) relevant to AI agent performance. Some example factors include whether the task involved a novel situation, was constrained by a finite resource, involved real-time coordination, or was sourced from a real-world context. We labeled RE-bench and HCAST tasks on the presence or absence of these 16 messiness factors, then summed these to obtain a “messiness score” ranging from 0 to 16. Factor definitions can be found in Appendix D.4.
However, trends in AI agent performance over time are similar for lower and higher messiness subsets of our tasks. For example, on sub hour tasks, success rates increased by 40 percentage points between Jan. 2023 and May 2025 in both high and low messiness splits (Figure 9). In particular, we find no evidence of either much slower performance trends, or a plateau, specific to our higher messiness subset.
Figure 10: We plot the excess success rate (the observed empirical task success rate, minus success rate we would predict using the task’s length, see Section 4.1) against messiness score for each task. As discussed in Section 6.2, there is a negative relationship between excess success rates and messiness.
@@ 6.3 SWE-bench Verified
To check whether we observe similar performance trends in time horizon on other benchmarks, we apply our methodology to SWE-bench Verified. SWE-bench Verified is an industry standard benchmark for evaluating language model performance on software engineering tasks [38]. All tasks in the SWE-bench Verified dataset were harvested from large open source repositories, like matplotlib or django, and then filtered to ensure they are automatically checkable and well-specified [20].
Figure 11: Performance of frontier AI models using reported SWE-bench Verified results (Section 6.3). We observe a similar exponential trend to Figure 1, albeit with a steeper slope.
Model time horizon computed from SWE-bench Verified tasks seem to follow an exponential trend from late 2023 through 2024. However, while the doubling time predicted by HCAST + SWAA + RE-bench using 2024 models is 104 days, the doubling time predicted by the SWE-bench Verified results is shorter—around 70 days.
SWE-bench Verified time annotations were based on the expected time it would take an “engineer who has had a few hours to familiarize themselves with the codebase” to solve the issue [20]. We found that annotator time estimates differentially underestimate how long our contract baseliners take to complete the easiest SWE-bench verified tasks. As a result, our time horizon estimates for SWE-bench Verified (which use annotator times) are likely to underestimate the time horizon of less capable models relative to contractor times, in turn shortening doubling times. For more details see Appendix D.3.
6.4 Internal PR experiments
We also ran GPT-4o, Claude 3.5 Sonnet (New), and o1 on five uncontaminated issues from an internal METR repository. Resolving these issues was real work performed by METR staff, so we might expect results on these tasks to better represent performance on real economically valuable tasks than a typical benchmark task.
We find that our contract baseliners take 5x-18x longer to resolve issues than repository maintainers. Additionally, AI agent performance on these issues is not inconsistent with AI agent success rate curves derived from HCAST, SWAA, and RE-Bench performance if contractor time-to-complete is used to measure the tasks length. However, it takes much longer for our contract baseliners to complete these tasks than repository maintainers. This suggests that time horizons may have better correspondence to the labor of a low-context human, rather than a high-context human. See Appendix B.2 for methodological details and more results.
@@ 7 Extrapolation
7.1 Extrapolating towards one-month-horizon AI
When forecasting when AI systems will be capable of autonomously generating large economic value and when they will be capable of catastrophic actions, it is necessary to choose a concrete threshold for horizon length. We chose one month (approximately 167 working hours for a fair comparison with humans, since humans cannot work 24/7) for two reasons. First, Ngo [31] writes that a 1-month AGI (defined as an AI that outperforms most knowledgeable humans who are given 1 month of work hours, i.e. 167 hours, to perform the task) would necessarily exceed human performance both at tasks including writing large software applications or founding startups (clearly economically valuable), and including novel scientific discoveries. [^16] Note that our forecasts concern AI with a 1-month horizon on software tasks, not 1-month AGI, because we evaluate models only on software and research tasks. Nevertheless, given past correlations in different areas of AI performance, 1-month (167 hours) time horizon AI may be significantly generally capable. Second, one month is around the period when new hires at a company begin to complete onboarding and generate economic value, [^17] Onboarding “can last from a few weeks to more than a year” [39], and employees often start generating economic value midway through the onboarding process. and so an AI with a horizon of 1 month could be capable of acquiring context like a human employee, allowing it to complete high-context as well as low-context tasks.
In this section, we attempt to forecast when AIs will reach a 50%-time horizon of 1 month, because it intuitively seems that a system capable of this length of task, even at 50% reliability, would be transformative for our society, including potentially being proficient in capabilities that could threaten society with catastrophic harm.
@@ 7.1 Extrapolating towards one-month-horizon AI
Although extrapolation is always imperfect, forecasting AI horizon lengths far into the future is much more sensitive to changes in doubling rate than to constant factors in horizon length. For example, a naive extrapolation based on o1’s time horizon of 39 minutes and a past doubling time of 218 days (3.2x/year) predicts that AIs will reach a 1-month 50% time horizon (roughly 8 doublings over o1) about 4.8 years after the release date of o1. A 2x increase in doubling time would delay the 1 month point by a further 4.8 years, but a 2x constant factor decrease in horizon length would only delay it by 0.6 years.
@@
Figure 12 shows a sensitivity analysis of our results over various sources of noise. “Bootstrap (tasks)” reflects our limited number of tasks and the fact that tasks of the same human duration vary in difficulty for models, while “Bootstap (runs)” derives from our limited number of runs; the combination of these two is shown as the confidence region in Figure 1. “Bootstrap (models)” represents the limited number of models in our dataset. “Weighting/regularization” represents methodological choices; this perturbation includes 10 hyperparameter combinations with and without task diversity weighting, and with a logistic regression regularization parameter between 0.01 and 0.2. “IID Baseline Noise” multiplies every task duration by a random factor based on the empirical distribution of baseline times. We combine these into an overall estimate of uncertainty in the past trend, and do the same for the 2024–2025 only trend (note that our confidence in this trend is much lower because there are only six models in this time span).
@@
Because our analysis puts low probability on the growth rate of time horizon on our tasks being much slower than one doubling every 8 months, the uncertainty in the extrapolated date of 1-month AI is fairly small (80% CI width about 2 years, central estimate late 2029). Bootstrapping over our task distribution and runs contributes most of the uncertainty, representing our limited task dataset and the variance in difficulty (for models) between tasks of the same human length, as well as run-to-run variance. If future progress instead follows the 2024–2025 trend, 1-month AI would arrive sooner, with half the probability in 2027 and early 2028.
@@
It is possible that systematic biases and alternate methodologies have a greater impact on forecasts, and we discuss these in Appendix D. In addition, due to the inherent difficulty of predicting the future, real forecasts will have larger error than this naive extrapolation, which we discuss below.
7.2 Difficulties in extrapolation
Most of our uncertainty about the future comes from (a) applicability to real tasks, partly discussed in Section 6, and (b) future changes in the time horizon growth rate.
7.2.1 Systematic differences between our tasks and real tasks
The tasks we use to benchmark AI capabilities are systematically different from real tasks. These differences could result in the trends we observe on our tasks not generalizing to real world tasks. For instance, all SWAA, HCAST, and RE-Bench differ from real-world tasks in the following ways:
  • No interaction with other agents None of our tasks involve interacting with other autonomous agents. Coordinating with, or competing with, other agents seems likely to increase task difficulty. For instance, by increasing the importance of strategic decision-making, real-time coordination, and predicting the actions of other complex agents.
  • Lax resource constraints None of our SWAA tasks, and few of our HCAST tasks saliently involve making efficient use of a limited resource—a common constraint in real-world tasks.
  • Unpunishing Similarly, very few of our tasks are punishing of single mistakes. [^18] With the exception of submitting an answer too early on HCAST tasks. This is in part to reduce the expected cost of collecting human baselines. Real world tasks can often be more punishing, for instance, when they involve competing against other agents. For instance, a single blunder in a chess game can greatly reduce the chance of winning the game.
  • Static environments Our tasks typically use environments that do not significantly change unless directly acted upon by the agent. In contrast, real tasks often occur in the context of a changing environment.
We attempted to measure how these systematic differences might affect AI agent performance in Section 6.2, by including the above properties as “messiness” factors. We found that though the absolute performance on “messier” tasks was lower, the trends in performance were similar to less messy tasks. Even so, these systematic differences cast doubt on whether the rapid performance improvements seen on our tasks (and other benchmarks like SWE-Bench Verified) will generalize to real world tasks.
Regardless of whether or not this trend generalizes to real world tasks, we believe our results to be significant. If our results do not generalize to real tasks, then benchmarks like HCAST and SWE-Bench Verified may be insufficient for forecasting AI capabilities on real tasks, and we may need more realistic benchmarks. On the other hand, if our results do generalize to real tasks, then extrapolating our trend predicts that AIs capable of automating a month of human software development will be made before 2032.
7.2.2 Future changes in time horizon trends
Here, we discuss three additional possible factors that could significantly change the time horizon growth rate: agency training, compute scaling, and automation of AI research and development.
Agency training
Compute scaling
Between the release of GPT-2 and today, the compute used to train the most impressive frontier language models has increased by at least a factor of 10,000x [23], with training compute usage doubling every 6–10 months [24]. More recently, models like o1 and o3 have began to use more compute at inference time. It is unclear whether there is sufficient capacity to expand either training or inference compute by many more orders of magnitude in the next 5 years. However, algorithmic improvements, which have historically decreased the compute requirements for a fixed performance level [40] [41], can substitute for compute limitations. We think that limits to compute scaling will slow the growth of AI agent time horizons somewhat, but be partially compensated by more investment into algorithmic improvement.
Automation of AI R&D

### v3

@@ 3.1 Computing time horizon
To convert agent run data to task completion time horizon, we first convert the agent performance on each task to a binary value (success or failure). Many tasks are naturally binary, including all SWAA tasks. Continuously scored tasks are binarized via a task-specific threshold chosen to represent human performance. For HCAST, the task-specific threshold is the same “target score” the human baseliner tries to achieve, which we also use to filter for successful runs. RE-Bench tasks have a fixed time rating of 8 hours, so the task-specific threshold is the average score of 7-9–hour human runs.
Figure 4: Success rates of all models on the test suite, showing the computation of time horizon as predicted 50% success rate time. The logistic fit is fairly good, though there is a jump in success rate between <1 minute SWAA tasks and >1 minute HCAST tasks.
@@
Once we have agent success rates and time ratings for each task, we fit time horizons by performing the following logistic regression: [^5] Compare to Item Response Theory (IRT) [3]. Like in IRT, we use logistic regression to find the task difficulty at which the agent has a 50% chance of success, but unlike IRT, we use difficulty ratings directly based on human baseline time rather than ratings learned from agent performance. Further details and comparison to standard IRT methods are provided in Appendix C.4.
$p_{\mathrm{success}}(\mathrm{agent},\mathrm{task})=\sigma((\log h_{\mathrm{agent}}-\log t_{\mathrm{task}})\cdot\beta_{\mathrm{agent}})$
@@
where $t_{\mathrm{task}}$ is the geometric mean time of successful human baselines, and $h_{\mathrm{agent}}$ and $\beta_{\mathrm{agent}}$ are learned parameters, with $h_{\mathrm{agent}}$ representing the 50% time horizon.
3.2 Time horizon vs. release date
In Figure 1, we plot the time horizons of each model against their release date—the date at which the lab first publicly announced the frontier model. [^6] In most cases, the release date is for each frontier AI model is the same model as the one we evaluated. However, GPT-3 (davinci) and GPT-3.5 (code-davinci-002 and text-davinci-002) are closed-source and no longer available through API access, but we use their release dates for the closest available models, which OpenAI advertises as having equivalent performance: GPT-3’s release date for davinci-002, and GPT-3.5’s date for gpt3.5-turbo-instruct. In addition, we linearly regress [^7] Specifically, we perform Ordinary Least Squares regression on $\log(\text{model\_horizon})=\alpha+\beta\cdot\text{release\_date}$. In Appendix H we discuss other curve-fitting methods, and conclude that the fit is not sensitive to various hyperparameters such as regularization, weighting of tasks, or WLS vs. OLS. $\log(\text{time horizon})$ against release date, finding that time horizon has doubled every 207 days with a 95% bootstrapped confidence interval 166-240 days (roughly $\pm 19\%$). Error bars are calculated via 10,000 samples from a three-level hierarchical bootstrap over task families, then tasks, then runs.
@@
Horizon length increases substantially over the entire time period from 2019 to early 2025. GPT-2 has a 50% time horizon of only 2 seconds, while o3 has a 110-minute time horizon and succeeds at several tasks over 4 hours. Also, o3 lies above the long-run trend (p = 0.006); this is robust to methodological ablations like using continuous scoring (Appendix H), which may imply the trend in 2024 and early 2025 is faster.
@@ 3.2.1 Time horizons at 50% success rate vs 80% success rate
To check whether our choice of 50% success rate affects the long-run trend, we also compute the time horizon at which AI agents succeed at tasks with 80% success rate, shown in Figure 17. The doubling time in 80% time horizon (204 days) is similar to the doubling time of 50% time horizon (207 days), within margin of error. However, models’ 80% time horizons are 4-6x shorter, suggesting that even models that sometimes succeed on difficult and diverse tasks cannot reliably perform tasks of moderate length. Due to the limited dataset, we cannot confidently measure time horizons at very high success rates (e.g. 95%)– see Section E.3.
3.3 Qualitative analysis
By examining transcripts from 31 randomly sampled unsuccessful GPT-4 1106 agent runs and 32 o1 runs, we observe that models seem to have improved greatly in terms of tool use capabilities, demonstrate a markedly greater ability to adapt to mistakes (as opposed to repeating unsuccessful actions), and perform much better at parts of tasks requiring logical reasoning or code generation. However, we noticed that AI agents still seem to struggle in intuitively “messier” environments—specifically, environments without clear feedback loops, or where the agent needs to proactively seek out relevant information.
Table 3.3 shows the categorized failures by model. We provide definitions of these categories, as well as examples of qualitative improvements and continuing limitations in Appendix D.
Table 2: Categorization of 31 failed runs by GPT-4 1106 and 32 failed runs by o1 (Section 3.3). Note that as o1 succeeds at more tasks, its failures correspond to more challenging tasks compared to GPT-4’s failures.
o 1 lcc Failure type GPT-4 1106 o1
@@ 4 External validity and robustness
We check whether the 2023–2025 trend without the SWAA dataset retrodicts the trend since 2019, and find that the trends agree. We also perform three experiments to check external validity.
First, we replicate our methods on SWE-bench Verified. We find that the exponential trend still holds (Figure 5), albeit with an even shorter doubling time, possibly because SWE-bench Verified difficulty annotations, meant to represent high-context maintainer time, may differentially underestimate how long human contractors take to perform easier SWE-bench tasks.
Second, we score the HCAST and RE-Bench tasks against a list of 16 “messiness” factors that aim to capture some systematic differences between the tasks studied here and hard“real-world” tasks. Some examples include whether the task is resource limited, novel, or involves a dynamic environment (Section H.6). Controlling for task length, we find models perform worse on tasks that have higher messiness scores. Notably, trends in AI agent performance over time are similar for the lower and higher messiness subsets of these tasks (see Figure 12). In particular, we find no evidence of plateaus in performance trends specific to the higher messiness subset.
Third, we measure AI agent performance on a small, uncontaminated set of our internal pull requests (PRs) (Section C.2). We find large differences in the speed at which different human groups complete the internal PR tasks, with contractors taking 5-18x longer to fix issues than repo maintainers. We also find that AI agent performance is worse than would be predicted by maintainer time-to-complete as a measure of task length, but is consistent with contractor time-to-complete,
SWE-bench Verified information is provided below. Details of the retrodiction and messiness analysis can be found in Appendix F, and details of the internal PRs experiment in C.2.
@@ 4.1 SWE-bench Verified
To check whether we observe similar performance trends on other benchmarks, we apply our methodology to SWE-bench Verified, an industry standard benchmark for evaluating language model performance on software engineering tasks [30]. All tasks in the SWE-bench Verified dataset were harvested from large open source repositories, like matplotlib or django, and then filtered to ensure they are automatically checkable and well-specified [19].
Figure 5:
Performance of frontier AI models using reported SWE-bench Verified results (Section 4.1). We observe a similar exponential trend to Figure 1, albeit with a steeper slope.
Model time horizon computed from SWE-bench Verified tasks seem to follow an exponential trend from late 2023 through 2024. However, while the doubling time predicted by HCAST + SWAA + RE-bench using 2024 models is 143 days, the doubling time on SWE-bench Verified results is shorter—around 70 days.
SWE-bench Verified time annotations were based on the expected time it would take an “engineer who has had a few hours to familiarize themselves with the codebase” to solve the issue [19]. We found that annotator time estimates differentially underestimate how long our contract baseliners take to complete the easiest SWE-bench verified tasks. As a result, our time horizon estimates for SWE-bench Verified (which use annotator times) are likely to underestimate the time horizon of less capable models relative to contractor times, in turn shortening doubling times. For more details see Appendix H.5.
@@ 5 Extrapolation
Extrapolating towards one-month-horizon AI
To forecast when AI systems may be capable of autonomously producing large economic value and potential catastrophic actions, we chose one month (approximately 167 working hours for a fair comparison with humans, since humans cannot work 24/7) for two reasons. First, Ngo [29] writes that a 1-month AGI (defined as an AI that outperforms most knowledgeable humans who are given 1 month of work hours, i.e. 167 hours, to perform the task) would necessarily exceed human performance at economically valuable endeavors like writing large software applications, founding startups, and making novel scientific discoveries. [^8] Note that our forecasts concern AI with a 1-month horizon on software tasks, not 1-month AGI, because we evaluate models only on software and research tasks. Nevertheless, given past correlations in different areas of AI performance, 1-month (167 hours) time horizon AI may be significantly generally capable. Second, one month is around the period when new hires at a company begin to complete onboarding and generate economic value, [^9] Onboarding “can last from a few weeks to more than a year” [52], and employees often start generating economic value midway through the onboarding process. and so an AI with a horizon of 1 month could be capable of acquiring context like a human employee, allowing it to complete high-context as well as low-context tasks. We cannot rule out that a system capable of 1-month tasks, even at 50% reliability, could be transformative for our society, potentially including the ability for catastrophic harm.
@@
Figure 6 shows a sensitivity analysis of our results over various sources of noise. “Bootstrap (tasks)”, “Bootstrap (runs)” and “Bootstrap (models)” represents the limited number of tasks, run-to-run variance, and limited models respectively. “Weighting/regularization” represents methodological choices; this perturbation includes 10 hyperparameter combinations with and without task diversity weighting, and with a logistic regression regularization parameter between 0.01 and 0.2. “IID Baseline Noise” multiplies every task duration by a random factor based on the empirical distribution of baseline times. Note our confidence in the 2024–2025 only trend is low because there are only seven frontier models in this time span.
@@
Because our analysis puts low probability on the growth rate of time horizon on our tasks being much slower than one doubling every 8 months, the uncertainty in the extrapolated date of 1-month AI is fairly small (80% CI width about 2 years, central estimate mid-2029). If future progress instead follows the 2024–2025 trend, 1-month AI would arrive sooner, with half the probability in 2027 and even late 2026.
@@
It is possible that systematic biases and alternate methodologies have a greater impact on forecasts, and we discuss these in Appendix H. In addition, due to the inherent difficulty of predicting the future, real forecasts will have larger error than this naive extrapolation, which we discuss below.
Task distribution limitations
We note several differences between this task suite and realistic tasks in Appendix B.2: in short, realistic tasks rarely have automatic scoring, and often require dealing with other agents, resource constraints, dynamic environments, and high reliability requirements. Such differences cast doubt on whether the rapid performance improvements seen on this task suite (and other benchmarks) will generalize. It may be that benchmarks like HCAST and SWE-Bench Verified are insufficient for forecasting real-world AI capabilities, and accurate forecasting will require deliberate construction of more realistic benchmarks without these limitations.
Future changes in time horizon trends
The time horizon trend may significantly change in the future due to factors including agency training, compute scaling, and automation of AI research and development. Agency training and eventual automation of AI R&D are more likely to speed up the trend, whereas compute scaling could slow it (but be partially offset by algorithmic improvements); a future artificial general intelligence (AGI) could technically have an infinite time horizon. See Appendix E for details.